When a car's starter is in use, it draws a large current. The car's lights draw much less current. As a certain car is starting, the current through the battery is 65.4 A and the potential difference across the battery terminals is 9.81 V. When only the car's lights are used, the current through the battery is 2.16 A and the terminal potential difference is 12.4 V. Find the battery's emf. Find the internal resistance.
Answer #1

Starter in use,

E = 9.81 + 65.4r

E - 65.4 r = 9.81 ....... (1)

Lights in use,

E = 12.4 + 2.16r

E - 2.16 r = 12.4 ....... (2)

Subtracting (1) - (2)

E = 12.49 V

r = 0.041 ohms

Comment in case any doubt please rate my answer ....